# Beyond Ensemble Averages: Leveraging Climate Model Ensembles for Subseasonal Forecasting

### v3

@@ Acknowledgements
Appendix A Regression results for NASA-GMAO and NCEP-CFSv2
@@ A.1 Regression
Precipitation regression results on the test data from NASA-GMAO are presented in Table 12. On this dataset, no learned method or method leveraging ensemble model forecasts significantly outperforms climatology. Note that the best $R^{2}$ value, associated with the climatology, is still negative; the low values for all methods indicate the difficulty of the forecasting problem.
Table 12: Test results for precipitation regression using NASA-GMAO dataset. LR refers to linear regression on all features including ensemble members, lagged data, land variables, and SSTs. Model stacking is performed on models that are learned on all features. Bold values indicate the best performance for each statistic. MSE is reported in mm.
@@
Figure 8 illustrates test performance of key methods on NASA-GMAO data with $R^{2}$ heatmaps over the U.S. Although the stacked model does not show the best performance in terms of mean $R^{2}$ score, it has more geographic regions with positive $R^{2}$ than any other method.
Figure 8: Test $R^{2}$ score heatmaps of baselines and learning-based methods for precipitation regression using the NASA-GMAO dataset. Positive values (blue) indicate better performance. See Section A.1 for details.
@@
Table 13: Test results for temperature regression using NASA-GMAO dataset. LR refers to linear regression on all features including ensemble members, lagged data, and land variables. Model stacking is performed on models that are learned on all features except SSTs. Bold values indicate the best performance for each statistic. MSE is reported in ${}^{\circ}C$.
@@
Figure 9: Test $R^{2}$ score heatmaps of baselines and learning-based methods for temperature regression using NASA-GMAO dataset. Positive values (blue) indicate better performance. See Section A.1 for details.
@@ A.2 Quantile regression
In Figure 10, we show heatmaps of quantile loss using of all locations in the U.S., where blue means smaller quantile loss and yellow means larger quantile loss. We observe that the learning-based models are outperforming the baselines, especially in Washington, California, Idaho, and near the Gulf of Mexico.
Figure 10: Test quantile loss heatmaps of baselines and learning-based methods for precipitation quantile regression using NCEP-CFSv2 dataset. Blue regions indicate smaller quantile loss. See Section 7.2 for details.
@@
Table 14 summarizes results for precipitation quantile regression using the NASA-GMAO ensemble. The models are the same as the models applied to the NCEP-CFSv2 ensemble. Our best model shows similar performance to that of the baselines. One possible reason is that the NASA-GMAO ensemble shows worse performance than the NCEP-CFSv2 ensemble empirically. Furthermore, according to the designs of ensemble members from both climate models, the NASA-GMAO ensemble has fewer ensemble members and therefore has less coverage on the distribution of precipitation than the NCEP-CFSv2 ensemble, so that our learned model has access to less information about the true distribution of precipitation.
@@
Figure 11: Test quantile loss heatmaps of baselines and learning-based methods for precipitation quantile regression using NASA-GMAO dataset. Blue regions indicate smaller quantile loss. See Section A.2 for details.
@@
Figure 12: Test quantile loss heatmaps of baselines and learning-based methods for temperature quantile regression using NASA-GMAO data. Blue regions indicate smaller quantile loss. See Section A.2 for details.
@@ B.1 Tercile classification of precipitation
In this case, the proposed learning-based methods are trained on the classification task directly. Predictions of baselines, such as the ensemble mean, are split into three classes according to the 33rd and 66th percentile values. Note that random forest and U-Net are trained for classification using all available features. We do not notice a significant difference in performance of logistic regression on the validation if the inputs are ensemble members only or ensemble members with side information. So, we use logistic regression on ensemble members only. The model stacking is applied to the logistic regression, U-Net and random forest outputs.
@@
Table 16: Test results for tercile classification of precipitation on different datasets. Accuracy in % is reported. Note that for this task, our models are trained for classification directly while baselines perform regression and threshold for predicted values is applied. For stacking, logistic regression, U-Net and RF outputs are used.
@@
Figure 13: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of precipitation using NCEP-CFSv2 dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Section B.1 for details.
The accuracy heatmaps over U. S. land are presented in the Figure 14 for NASA-GMAO dataset. The plots corresponding to the learning-based methods show the best results, the ensemble mean’s figure has the most red regions.
Figure 14: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of precipitation using NASA-GMAO dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Section B.1 for details.
@@ B.2 Tercile classification of temperature
Figure 15 shows accuracy heatmaps over the U.S. for different methods using NCEP-CFSv2 data. The stacked model shows the best performance across spatial locations. For example, the ensemble mean does not show great performance in the South East and Middle Atlantic regions, while learning based methods demonstrate much stronger predictive skill in these areas. However, there are still some areas, such as Texas or South West region, with red pixels for all methods.
Figure 15: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of temperature using NCEP-CFSv2 dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Section B.2 for details.
@@
Figure 16: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of temperature using NASA-GMAO dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess.See Section B.2 for details.
@@ Appendix C Extended Discussions
In this section, we present more detailed results for Section 8. We also provide additional experiments on the temperature forecasting analysis, and experiments on training set sizes and bootstrap.
C.1 Ensemble formulation
In Section 8.1, we show a set of experiments with different usage of the ensemble. Additionally, we conduct an experiment designed to test whether it is important to keep track of which ensemble member made each prediction or whether it is the distribution of predictions that is important. The modeling approach for the former would be to feed in ensemble member 1’s forecast as the first feature, ensemble member 2’s forecast as the second feature, etc. The modeling approach under the distributional hypothesis is to make the smallest prediction be the first feature, the second-smallest prediction be the second feature, and so on – i.e., we sort the ensemble forecasts for each location separately. Note that this entails treating the ensemble members symmetrically: the model would give the same prediction if ensemble member 1 predicted $a$ and ensemble member 2 predicted $b$ or if ensemble member 1 predicted $b$ and ensemble member 2 predicted $a$.In statistical parlance, this is passing in the order statistics of the forecasts as the features, rather than their original ordering. Note that for NCEP-CFSv2, ensemble forecasts are originally ordered according to the time at which their initial conditions are set [15].
C.2 PE v.s. latitude/longitude values or no location information
In this section, we elaborate on our experiment with different usage of location information. Similar to Section 85.4, we train and test our models with three different settings: using no location information, using latitude/longitude values or using positional encodings. For the stacked model, we first train the LR, RF and U-Net using these different settings, before training the stacked model using the corresponding LR, RF and U-Net outputs. Table 18 summarizes test performance for precipitation regression with these three settings of using location information. For linear regression, we observe that having latitude/longitude values or adding PE features does not improve its performance. One interesting result is that adding PE for the LR degraded its performance, which may due to that the PE features are non-linear transformations of latitude and longitude values, which is hard to fit with a linear model.
For the RF, U-Net and the stacked model, we observe that adding PE improves their performance with significance, i.e. having at least one standard error smaller MSE. For the U-Net, using latitude/longitude values yields worse overall performance compared to using no location information. Figure 17 shows the test $R^{2}$ heatmaps for LR, U-Net, RF and stacked model under these three settings for precipitation regression. We observe that adding PE features not only improves performance for the U-Net and RF, the stacked model’s performance also improves from having better predictions from the U-Net and RF.
@@ C.1 Ensemble formulation
Table 19 shows the test performances of LR, U-Net, RF and stacked model on temperature regression. Similar to precipitation regression, adding latitude/longitude values or adding PE does not help the LR, but we observe significant performance improvement when adding positional encodings features for the U-Net, RF and the stacked model. Figure 18 then shows the test $R^{2}$ heatmaps for temperature regression. We can see that adding PE to the U-Net, RF and stacked model improves forecast performance, especially in regions like Arizona, New Mexico and Texas.
@@
C.3 Bootstrap experiments
To evaluate the stability of our machine learning models with small sample sizes, we perform the following bootstrap experiments: We take bootstrap samples of size 200 from our training set and retrain our U-Net, RF and LR. Then we evaluate these models on the test set. We repeat this process for 50 times and show the results in Fig. 19. We observe from the plots that the U-Net performs consistently better than the LR in precipitation regression, but not for temperature regression. This result is consistent with what we showed in Table 3 and Table 4.
@@ C.3 Bootstrap experiments
C.4 Precipitation forecast example
@@
Figure 20: An illustration of precipitation predictions ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{y}_{t,l}^{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}anomaly}}}}$ (in mm) of different methods for February 2016 (in test period). (a) True precipitation. (b) LR on ensemble members. (c) Climatology. (d) LR on all features. (e) Ensemble mean. (f) U-Net on all features. (g) Example single ensemble member. (h) Random forest on all features. (i) Stacked model. See Section C.4 for details.
Figure 21 demonstrates differences between the ground truth and different model predictions. In this figure, the color white is associated with the smallest errors, while red pixels indicate overestimating precipitation and blue pixels indicate underestimating precipitation. The individual ensemble member in Figure 21(e) exhibits dark red regions across the West, while the ensemble mean in Figure 21(e) shows better performance in this area. The colors are more muted for the stacked model in Figure 21(h). The climatology in Figure 21(a) has the most neutral areas. However, its MSE is slightly higher than the stacked model’s MSE. In general, all methods, including linear regression (b, d), U-Net (f) and random forest (g), tend to underpredict precipitation in the Southeast, Mid-Atlantic and North Atlantic and predict higher precipitation levels in the West.
Figure 21: An illustration of differences $y_{t,l}^{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}anomaly}}}-\hat{y}_{t,l}^{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}anomaly}}}$ in precipitation predictions in mm of different methods for February 2016 (in test period). Red pixels indicate areas where a forecasting method predicts higher precipitation levels compared to the ground truth, blue pixels indicate an underestimation of the precipitation, and white pixels correspond to a precise prediction. See Section C.4 for details.
C.5 Temperature forecasting analysis
Figure 3 shows regions in Texas and Florida where the ensemble mean and linear regression performance is poor, while a random forest achieves far superior performance. We conduct an analysis of forecasts of the ensemble mean, linear regression, and random forests in these regions together with a region in Wisconsin where all methods show good performance. Figure 22 indicates these regions and Table 20 summarizes performance of different methods in these regions: the ensemble mean prediction quality dramatically drops between the validation and test periods in Texas and Florida, which is not the case for the random forest.
Figure 22: Regions where the temperature forecast is analyzed. See Section C.5 for details.
@@ C.5 Temperature forecasting analysis
Why does RF perform so much better than simpler methods in some regions? One possibility is that the RF is a nonlinear model capable of more complex predictions. However, if that were the only cause of the discrepancy in performance, then we would expect that the RF would be better not only on the test period, but on the validation period as well. Table 20 does not support this argument; it shows that the ensemble mean and linear regression have comparable if not superior performance to the random forest during the validation period. A second hypothesis is that the distribution of temperature is different during the test period than during the training and validation periods. This hypothesis is plausible for two reasons: (1) climate change, and (2) the training and validation data use hindcast ensembles while the test data uses forecast ensembles. To investigate this hypothesis, in Figure 23 we plot the true temperature and ensemble mean in the training, validation, and test periods for the three geographic regions. The discrepancy between the true temperatures and ensemble means in the test period is generally greater than during the training and validation periods in Texas and Florida (though not in Wisconsin, a region in which validation and test performance are comparable for all methods). This lends support to the hypothesis that hindcast and forecast ensembles exhibit distribution drift, and the superior performance of the RF during the test period may be due to a greater robustness to that distribution drift.
@@
The hindcast and forecast ensembles may have different predictive accuracies because the hindcast ensembles have been debiased to fit past observations – a procedure not possible for forecast data. To explore the potential impact of debiasing, Figure 23 shows the “oracle debiased ensemble mean”, which is computed by using the test data to estimate the forecast ensemble bias and subtract it from the ensemble mean. This procedure, which would not be possible in practice and is used only to probe distribution drift ensemble bias, yields smaller discrepancies between the true data and the (oracle debiased) ensemble mean than the discrepancies between the true data and the original (biased) ensemble mean. Specifically, the oracle ensemble member achieves -0.20 mean $R^{2}$ score (TX) and -0.28 mean $R^{2}$ score (FL) vs. -1.55 $R^{2}$ (TX) and -0.87 $R^{2}$ (FL) of the original forecast ensemble mean. The errors during the test period are generally larger than during the train and validation period, even after debiasing the ensemble members using future data. This effect may be attributed both to (a) the nonstationarity of the climate (note that there are more extreme values during the test period than during the training and validation periods, particularly in Texas and Florida) and (b) the fact that in the train and validation periods, we use hindcast ensemble members, whereas in the test period, we use forecast ensemble members.
@@
Figure 23: Temperature predictions in ${}^{\circ}C$ of different methods at Texas, Florida, and Wisconsin regions. Black lines correspond to train/val and val/test splits; train and validation correspond to the hindcast regime of the ensemble, while test corresponds to the forecast regime. See Section C.5 for details.
Appendix D Architecture Details

### v4

@@ Acknowledgments
Appendix A Regression results for for NASA-GMAO Dataset
@@ A.1 Regression
Precipitation regression results on the test data from NASA-GMAO are presented in Table 12. On this dataset, no learned method or method leveraging ensemble model forecasts significantly outperforms climatology. Note that the best $R^{2}$ value associated with the climatology is still negative; the low values for all methods indicate the difficulty of the forecasting problem.
Table 12: Test results for precipitation regression using NASA-GMAO dataset. LR refers to linear regression on all features including ensemble members, lagged data, land variables, and SSTs. Model stacking is performed on models that are learned on all features. Bold values indicate the best performance for each statistic. MSE is reported in squared mm.
@@
Figure 8 illustrates the test performance of key methods on NASA-GMAO data with $R^{2}$ heatmaps over the U.S. Although the stacked model does not show the best performance in terms of mean $R^{2}$ score, it has more geographic regions with positive $R^{2}$ than any other method.
Figure 8: Test $R^{2}$ score heatmaps of baselines and learning-based methods for precipitation regression using the NASA-GMAO dataset. Positive values (blue) indicate better performance. See Appendix A.1 for details.
@@
Table 13: Test results for temperature regression using NASA-GMAO dataset. LR refers to linear regression on all features including ensemble members, lagged data, and land variables. Model stacking is performed on models that are learned on all features except SSTs. Bold values indicate the best performance for each statistic. MSE is reported in squared ${}^{\circ}C$.
@@
Figure 9: Test $R^{2}$ score heatmaps of baselines and learning-based methods for temperature regression using NASA-GMAO dataset. Positive values (blue) indicate better performance. See Appendix A.1 for details.
@@ A.2 Quantile regression
In Figure 10, we show heatmaps of quantile loss using all locations in the U.S., where blue means smaller quantile loss and yellow means larger quantile loss. We observe that the learning-based models outperform the baselines, especially in Washington, California, Idaho, and near the Gulf of Mexico.
Figure 10: Test quantile loss heatmaps of baselines and learning-based methods for precipitation quantile regression using NCEP-CFSv2 dataset. Blue regions indicate smaller quantile loss. See Appendix 7.2 for details.
@@
Table 14 summarizes results for precipitation quantile regression using the NASA-GMAO ensemble. The models are the same as the models applied to the NCEP-CFSv2 ensemble. Our best model shows similar performance to that of the baselines. One possible reason is that the NASA-GMAO ensemble shows worse performance than the NCEP-CFSv2 ensemble empirically. Furthermore, according to the designs of ensemble members from both climate models, the NASA-GMAO ensemble has fewer ensemble members and, therefore, has less coverage on the distribution of precipitation than the NCEP-CFSv2 ensemble, so our learned model has access to less information about the true distribution of precipitation.
@@
Figure 11: Test quantile loss heatmaps of baselines and learning-based methods for precipitation quantile regression using NASA-GMAO dataset. Blue regions indicate smaller quantile loss. See Appendix A.2 for details.
@@
Figure 12: Test quantile loss heatmaps of baselines and learning-based methods for temperature quantile regression using NASA-GMAO dataset. Blue regions indicate smaller quantile loss. See Appendix A.2 for details.
@@ B.1 Tercile classification of precipitation
In this case, the proposed learning-based methods are directly trained on the classification task. Predictions of baselines, such as the ensemble mean, are split into three classes according to the 33rd and 66th percentile values. Note that random forest and U-Net are trained for classification using all available features. We do not notice a significant difference in the performance of logistic regression on the validation if the inputs are ensemble members only or ensemble members with side information. So, we use logistic regression on ensemble members only. The model stacking is applied to the logistic regression, U-Net, and random forest outputs.
@@
Table 16: Test results for tercile classification of precipitation on different datasets. Accuracy in % is reported. Note that for this task, our models are trained for classification directly while baselines perform regression, and a threshold for predicted values is applied. For stacking, logistic regression, U-Net and RF outputs are used.
@@
Figure 13: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of precipitation using NCEP-CFSv2 dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Appendix B.1 for details.
The accuracy heatmaps over the U. S. land are presented in Figure 14 for the NASA-GMAO dataset. The plots corresponding to the learning-based methods show the best results, the ensemble mean’s figure has the most red regions.
Figure 14: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of precipitation using NASA-GMAO dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Appendix B.1 for details.
@@ B.2 Tercile classification of temperature
Figure 15 shows accuracy heatmaps over the U.S. for different methods using NCEP-CFSv2 data. The stacked model shows the best performance across spatial locations. For example, the ensemble mean does not show great performance in the Southeast and Middle Atlantic regions, while learning-based methods demonstrate much stronger predictive skills in these areas. However, there are still some areas, such as Texas or South West region, with red pixels for all methods.
Figure 15: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of temperature using NCEP-CFSv2 dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Appendix B.2 for details.
@@
Figure 16: Test accuracy heatmaps of baselines and learning-based methods for tercile classification of temperature using NASA-GMAO dataset. The accuracy colorbar is recentered to be white at $\frac{1}{3}$, what corresponds to a random guess score. Blue pixels indicate better performance, while red pixels correspond to performance that is worse than a random guess. See Appendix B.2 for details.
@@ Appendix C Extended Discussions
In this section, we present more detailed results for Section 8.1 and Section 5.4. We also provide additional experiments on the temperature forecasting analysis, and experiments on training set sizes and bootstrap.
C.1 PE v.s. latitude/longitude values or no location information
In this section, we elaborate on our experiment with different uses of location information. Similar to Section 5.4, we train and test our models with three different settings: using no location information, using latitude/longitude values, or using positional encodings. For the stacked model, we first train the LR, RF, and U-Net using these different settings before training the stacked model using the corresponding LR, RF, and U-Net outputs. Table 18 summarizes test performance for precipitation regression with these three settings of using location information. For linear regression, we observe that having latitude/longitude values or adding PE features does not improve its performance. One interesting result is that adding PE for the LR degraded its performance, which may be due to the fact that the PE features are non-linear transformations of latitude and longitude values, which is hard to fit with a linear model.
For the RF, U-Net, and the stacked model, we observe that adding PE improves their performance with significance, i.e. having at least one standard error smaller MSE. For the U-Net, using latitude/longitude values yields worse overall performance compared to using no location information. Figure 17 shows the test $R^{2}$ heatmaps for LR, U-Net, RF, and stacked model under these three settings for precipitation regression. We observe that adding PE features not only improves performance for the U-Net and RF, the stacked model’s performance also improves from having better predictions from the U-Net and RF.
@@
Table 19 shows the test performances of LR, U-Net, RF and stacked model on temperature regression. Similar to precipitation regression, adding latitude/longitude values or adding PE does not help the LR, but we observe significant performance improvement when adding positional encoding features for the U-Net, RF, and the stacked model. Figure 18 then shows the test $R^{2}$ heatmaps for temperature regression. We can see that adding PE to the U-Net, RF, and stacked model improves forecast performance, especially in regions like Arizona, New Mexico, and Texas.
@@
C.2 Bootstrap experiments
To evaluate the stability of our machine learning models with small sample sizes, we perform the following bootstrap experiments: We take bootstrap samples of size 200 from our training set and retrain our U-Net, RF, and LR. Then we evaluate these models on the test set. We repeat this process 50 times and show the results in Fig. 19. We observe from the plots that the U-Net performs consistently better than the LR in precipitation regression but not for temperature regression. This result is consistent with what we showed in Table 3 and Table 4.
@@ C.2 Bootstrap experiments
C.3 Precipitation forecast example
@@
Figure 20: An illustration of precipitation predictions ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{y}_{t,l}^{\text{anomaly}}}$ ( in mm) of different methods for February 2016 (in test period). (a) True precipitation. (b) LR on ensemble members. (c) Climatology. (d) LR on all features. (e) Ensemble mean. (f) U-Net on all features. (g) Example single ensemble member. (h) Random forest on all features. (i) Stacked model. See Appendix C.3 for details.
Figure 21 demonstrates differences between the ground truth and different model predictions. In this figure, the color white is associated with the smallest errors, while red pixels indicate overestimating precipitation and blue pixels indicate underestimating precipitation. The individual ensemble member in Figure 21(e) exhibits dark red regions across the West, while the ensemble mean in Figure 21(e) shows better performance in this area. The colors are more muted for the stacked model in Figure 21(h). The climatology in Figure 21(a) has the most neutral areas. However, its MSE is slightly higher than the stacked model’s MSE. In general, all methods, including linear regression (b, d), U-Net (f), and random forest (g), tend to underpredict precipitation in the Southeast, Mid-Atlantic, and North Atlantic and predict higher precipitation levels in the West.
Figure 21: An illustration of differences $y_{t,l}^{\text{anomaly}}-\hat{y}_{t,l}^{\text{anomaly}}$ in precipitation predictions in mm of different methods for February 2016 (in test period). Red pixels indicate areas where a forecasting method predicts higher precipitation levels compared to the ground truth, blue pixels indicate an underestimation of the precipitation, and white pixels correspond to a precise prediction. See Appendix C.3 for details.
C.4 Temperature forecasting analysis
Figure 3 shows regions in Texas and Florida where the ensemble mean and linear regression performance is poor, while a random forest achieves far superior performance. We conduct an analysis of forecasts of the ensemble mean, linear regression, and random forests in these regions together with a region in Wisconsin where all methods show good performance. Figure 22 indicates these regions and Table 20 summarizes the performance of different methods in these regions: the ensemble mean prediction quality dramatically drops between the validation and test periods in Texas and Florida, which is not the case for the random forest.
Figure 22: Regions where the temperature forecast is analyzed. See Appendix C.4 for details.
@@ C.4 Temperature forecasting analysis
Why does RF perform so much better than simpler methods in some regions? One possibility is that the RF is a nonlinear model capable of more complex predictions. However, if that were the only cause of the discrepancy in performance, then we would expect that the RF would be better not only during the test period, but during the validation period as well. Table 20 does not support this argument; it shows that the ensemble mean and linear regression have comparable, if not superior, performance to the random forest during the validation period. A second hypothesis is that the distribution of temperature is different during the test period than during the training and validation periods. This hypothesis is plausible for two reasons: (1) climate change, and (2) the training and validation data use hindcast ensembles while the test data uses forecast ensembles. To investigate this hypothesis, in Figure 23 we plot the true temperature and ensemble mean in the training, validation, and test periods for the three geographic regions. The discrepancy between the true temperatures and ensemble means in the test period is generally greater than during the training and validation periods in Texas and Florida (though not in Wisconsin, a region where validation and test performance are comparable for all methods). This lends support to the hypothesis that hindcast and forecast ensembles exhibit distribution drift, and the superior performance of the RF during the test period may be due to a greater robustness to that distribution drift.
@@
The hindcast and forecast ensembles may have different predictive accuracies because the hindcast ensembles have been debiased to fit past observations – a procedure not possible for forecast data. To explore the potential impact of debiasing, Figure 23 shows the “oracle debiased ensemble mean”, which is computed by using the test data to estimate the forecast ensemble bias and subtracting it from the ensemble mean. This procedure, which would not be possible in practice and is used only to probe distribution drift ensemble bias, yields smaller discrepancies between the true data and the (oracle debiased) ensemble mean than the discrepancies between the true data and the original (biased) ensemble mean. Specifically, the oracle ensemble member achieves -0.20 mean $R^{2}$ score (TX) and -0.28 mean $R^{2}$ score (FL) vs. -1.55 $R^{2}$ (TX) and -0.87 $R^{2}$ (FL) of the original forecast ensemble mean. The errors during the test period are generally larger than during the train and validation period, even after debiasing the ensemble members using future data. This effect may be attributed both to (a) the nonstationarity of the climate (note that there are more extreme values during the test period than during the training and validation periods, particularly in Texas and Florida) and (b) the fact that in the train and validation periods, we use hindcast ensemble members, whereas in the test period, we use forecast ensemble members.
@@
Figure 23: Temperature predictions in ${}^{\circ}C$ of different methods at Texas, Florida, and Wisconsin regions. Black lines correspond to train/val and val/test splits; train and validation correspond to the hindcast regime of the ensemble, while test corresponds to the forecast regime. See Appendix C.4 for details.
Appendix D Machine learning architectures